Find and correct the errors in the following mathematical statements.
x(3x+2) = 3x^2 +2

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

LHS = x(3x+2) = 3x^2+2x ≠3x^2+2 = RHS
The correct solution is x(3x+2) = 3x^2+2x

Thanks